OpenStudy (anonymous):

What are the foci and ellipse given by the equation? 225x^2 + 144y^2=32,400

OpenStudy (anonymous):

make it look like \[\frac{a^2}{a^2}+\frac{y^2}{b^2}=1\] by dividing both sides by \(32,400\) then find \(c\) via \(c^2=a^2-b^2\)

OpenStudy (anonymous):

Can you help me find c? Do i plug in 225 and 144 for the equation? Thank you for your help :)